OpenStudy (lucaz):

plot the function f(x)=x^3/3 - x

OpenStudy (anonymous):

Use this: https://www.desmos.com/calculator

OpenStudy (lucaz):

what if I were on a test?

OpenStudy (anonymous):

the solution is approximately 1.5874010519682 which rounds to 1.6

OpenStudy (lucaz):

solution? I need the graph of the function, how to do the sketch

OpenStudy (anonymous):

Use the website then copy & paste your equation in the left-hand box.

OpenStudy (anonymous):

Well the line is on the x-axis.

OpenStudy (lucaz):

I know these sites, but I want to know how to do it without a tool

OpenStudy (anonymous):

Are you in Calculus? Calculus makes the graphing of this particular equation a bit easier.

OpenStudy (lucaz):

no

OpenStudy (anonymous):

Then I'd start by first figuring out where the zeros are. That is, where the function will cross the x-axis. To do that, set the equation = to 0 and solve x^3/3 - x = 0 x(x^2 -1) = 0 x(x +1)(x-1)=0 x = 0 or x = -1 or x = 1............mark dots on the x-axis at these points

OpenStudy (lucaz):

yeah, i did this

OpenStudy (lucaz):

the problem was to fin min/max between the zeros

OpenStudy (anonymous):

ah...ok! you want to find f(-1/2) and f(1/2) . That will tell us where the graph will reach its minimums and maximums between those zeros.

OpenStudy (lucaz):

exactly in the middle of -1 and 0, 0 and 1?

OpenStudy (anonymous):

I get f(-1/2) = 11/24 and f(1/2) = -11/24 So the graph will look like this